# Neutrinos, B-L Symmetry and the Dark Dimension

### v1

@@ 4 Concluding thoughts on neutrino masses and GUTs
It is reassuring that these features and intriguing correlations come out without fine tuning in what is perhaps the simplest embedding of a B-L bulk gauge field with a single extra dimension, once the Dark Dimension scaling $m_{KK}\sim\Lambda^{1/4}$ is imposed. In other words, it is precisely the lack of additional fine-tunings (both in the value of the Yukawa coupling and in the relation between the vevs of the Higgs and bulk field $\Phi$) that explains the intriguing coincidence between neutrino masses and dark energy. This simplicity, which has appeared in other aspects of the Dark Dimension scenario, suggests that it may be on the right track.
@@
Another thought is that, although not strictly required by the model, it is natural to put the SM brane at the end of an interval (see e.g. Schwarz:2024tet). In this case, there will be another boundary and, consequently, another brane at the far side of the Dark Dimension. This may or may not have gapless degrees of freedom (think of e.g. the T-dual of type I string theory on a circle), but its degrees of freedom may couple as B-L in a similar way to the SM brane. One should keep in mind this possibility as the B-L bulk scenario is explored further.
@@ Appendix A Anomalies for the bulk theory
In Section 3.2, we introduced a five-dimensional bulk theory with Pin+ reflections coupled to $U(1)$ gauge fields. To determine if there are any anomaly cancellation constraints on this fermion spectrum we must compute the bordism group $\Omega_{6}^{\text{Pin}^{+}}(BU(1))$. This can be done via the Atiyah-Hirzebruch spectral sequence (see e.g. Garcia-Etxebarria:2018ajm). The second page of this sequence is displayed in Figure 2.
@@
At degree six, there are three possible entries, and just two potentially relevant differentials. The differential $\beta$ has as target the only class in dimension 5, but this class (corresponding to $KB\times S^{1}\times S^{2}$ with one unit of $U(1)$ flux along the $S^{2}$) vanishes: just take the $S^{1}\times S^{2}$ factor and view it as a class in $\Omega_{3}^{\text{Spin}}(BU(1))=0$. This means the differential is nontrivial, and the target class in dimension 6 is also killed. The source of the differential $\alpha$ is $KB\times S^{1}\times S^{2}$, and is trivial for the same reason. On the other hand, the class in $(6,2)$, which could be the source of a differential to the source of $\alpha$, is represented by $\mathbb{C}^{3}$ with the canonical $U(1)$ bundle on top, which is non-trivial since it is detected by the 7d Dirac $\eta$ invariant. As a result, $\alpha$ is also non-trivial, and we are looking at an extension of a single $\mathbb{Z}_{2}$ factor by $\mathbb{Z}_{8}$. The $\mathbb{Z}_{2}$ is represented by $\mathbb{CP}^{3}$ with the canonical $U(1)$ bundle on top, and detected by the bordism invariant $\int c_{1}^{3}\,\text{mod}\,2$. So

### v2

@@ 4 Concluding thoughts on neutrino masses and GUTs
It is reassuring that these features and intriguing correlations come out without fine tuning in what is perhaps the simplest embedding of a B-L bulk gauge field with a single extra dimension, once the Dark Dimension scaling $m_{KK}\sim\Lambda^{1/4}$ is imposed. In other words, it is precisely the lack of additional fine-tunings (both in the value of the Yukawa coupling and in the relation between the vevs of the Higgs and bulk field $\Phi$) that explains the intriguing coincidence between neutrino masses and dark energy. This simplicity, which has appeared in other aspects of the Dark Dimension scenario, suggests that it may be on the right track. In this line, it would be interesting to test this scenario in the context of recent searches for extra-dimensional sterile neutrinos Elaçmaz et al. (2025); de Giorgi et al. (2025) [^6] We believe some of the Majorana results of de Giorgi et al. (2025) and related references likely need revision, since the prediction in a simple dimensional reduction of a 4d KK mode much lighter than the 5d bulk fermion mass violates Lorentz invariance..
@@
Another thought is that, although not strictly required by the model, it is natural to put the SM brane at the end of an interval (see e.g. Schwarz (2024)). In this case, there will be another boundary and, consequently, another brane at the far side of the Dark Dimension. This may or may not have gapless degrees of freedom (think of e.g. the T-dual of type I string theory on a circle), but its degrees of freedom may couple as B-L in a similar way to the SM brane. One should keep in mind this possibility as the B-L bulk scenario is explored further.
@@ Appendix A Anomalies for the bulk theory
In Section 3.2, we introduced a five-dimensional bulk theory with Pin+ reflections coupled to $U(1)$ gauge fields. To determine if there are any anomaly cancellation constraints on this fermion spectrum we must compute the bordism group $\Omega_{6}^{\text{Pin}^{+}}(BU(1))$. This can be done via the Atiyah-Hirzebruch spectral sequence (see e.g. García-Etxebarria and Montero (2019)). The second page of this sequence is displayed in Figure 2.
@@
At degree six, there are three possible entries, and just two potentially relevant differentials. The differential $\beta$ has as target the only class in dimension 5, but this class (corresponding to $KB\times S^{1}\times S^{2}$ with one unit of $U(1)$ flux along the $S^{2}$) vanishes: just take the $S^{1}\times S^{2}$ factor and view it as a class in $\Omega_{3}^{\text{Spin}}(BU(1))=0$. This means the differential is nontrivial, and the source class in dimension 6 is also killed. The source of the differential $\alpha$ is $KB\times S^{2}\times S^{2}\times S^{1}$, and is trivial for a similar reason. On the other hand, the class in $(6,2)$, which could be the source of a differential to the source of $\alpha$, is represented by $\mathbb{CP}^{3}\times KB$ with the canonical $U(1)$ bundle on the $\mathbb{CP}^{3}$ factor, which is non-trivial since it is detected by the 8d $\eta$ invariant of a Dirac fermion of charge 2 (so that the Dirac index is 1). As a result, $\alpha$ is also non-trivial, and we are looking at an extension of a single $\mathbb{Z}_{2}$ factor by $\mathbb{Z}_{8}$. The $\mathbb{Z}_{2}$ is represented by $\mathbb{CP}^{3}$ with the canonical $U(1)$ bundle on top, and detected by the bordism invariant $\int c_{1}^{3}\,\text{mod}\,2$. So